QUADRATIC MINIMA AND MODULAR FORMS

BARRY BRENT

Abstract. We give upper bounds on the size of the gap between the constant term and the next nonzero Fourier coefficient of an entire modular form of given weight for Γ_{0}(2). Numerical evidence indicates that a sharper bound holds for the weights h ≡ 2 (mod 4). We derive upper bounds for the minimum positive integer represented by level-two even positive-definite quadratic forms. Our data suggest that, for certain meromorphic modular forms and p = 2, 3, the p-order of the constant term is related to the base-p expansion of the order of the pole at infinity.